IMPORTANT Questions 1 & 2 ask for cash flows only, no present values. They are a critical part of the problem, but since the116 Mon 2: Tixta Smrt Prom Table 1: Present Value of $1.00 perind 1129 13% 14% 5% 6% 7% 0.980 1 0.971 0.990 06 10052 0.962 0.
Answer #1

Answer of 0 year 2 cash flow using curent system Cash flow = cost do inspection + (number detected 1 x cost to fix) + (numberRefund value = Defektive units shipped x lost to replace = (406-330) x (1.2 x 195) = 70 x 234 =$16380 Total cash flow = -6100Present valul factor@tt cash flows 238,000 238.000 0.943 0.890 238.000 0.840 238,000 0.792 238,000 0.747 Net Dresent value =Nct present value = cost of replacement present value dr annuity = - $392643 – 639000 = -1031643 L inet present valere - - $1